OpenStudy (anonymous):

can someone solve x^2+4x=6?

OpenStudy (anonymous):

what dux the ^ mean?

myininaya (myininaya):

x^2+4x+4=6+4 (x+2)^2=10 x+2=+ or -sqrt{10} x=-2 + or -sqrt{10}

OpenStudy (nikita2):

http://www.wolframalpha.com/input/?i=x^2%2B4x%3D6

OpenStudy (anonymous):

@nakealro it means squared

OpenStudy (anonymous):

oooo thankz

OpenStudy (anonymous):

x^2+4x-6=0 x=2 x-2